OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above? y = - x + 30 y = 3x + 6 y = - x + 6 y = x + 30

OpenStudy (anonymous):

c?

OpenStudy (texaschic101):

3x + 5y = 30 5y = -3x + 30 y = -3/5x + 6.....its not there

OpenStudy (anonymous):

\(\Huge\color{red}{ \bf }\)

OpenStudy (anonymous):

The area of a parallelogram is A = bh. Solve for h (height) in this formula.